OpenStudy (anonymous):

How do you solve this equation 4nsquared-81=0 using the quadratic formula?

jimthompson5910 (jim_thompson5910):

In the case of \[\Large 4n^2-81=0\], its in the form \[\Large an^2+bn+c=0\] where a = 4, b = 0 and c = -81 Now use the quadratic formula to solve \[\Large n = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large n = \frac{-(0)\pm\sqrt{(0)^2-4(4)(-81)}}{2(4)}\] \[\Large n = \frac{0\pm\sqrt{0-(-1296)}}{8}\] \[\Large n = \frac{\pm\sqrt{1296}}{8}\] \[\Large n = \frac{\sqrt{1296}}{8} \ \text{or} \ n = \frac{-\sqrt{1296}}{8}\] \[\Large n = \frac{36}{8} \ \text{or} \ n = \frac{-36}{8}\] \[\Large n = \frac{9}{2} \ \text{or} \ n = \frac{-9}{2}\] So the solutions are \[\Large n = \frac{9}{2} \ \text{or} \ n = \frac{-9}{2}\]